OpenStudy (anonymous):

What is the slope of the line between (–4, 4) and (–1, –2)?

OpenStudy (anonymous):

3/6=1/2

OpenStudy (anonymous):

Knowing points (-4, 4) and (-1, -2), the slope, m, is given by: m = [(y2) - (y1)]/[(x2) - (x1)] = [(-2) - (4)]/[(-1) - (-4)] = -6/3 = -2

OpenStudy (anonymous):

thank you